At a unit price of $768, the quantity demanded of a certain commodity is 73 pounds. If the unit price increases to $844, the quantity demanded decreases by 19 pounds. Find the demand equation (assuming it is linear) where p is the unit price and x is the quantity demanded for this commodity in pounds.
p = ?



At what price are no consumers willing to buy this commodity?
$  

According to the above model, how many pounds of this commodity would consumers take if it was free?
  pounds
Answer #1

If x is the quantity demanded in posride and p is the unit price for a unil Doice of Pi=8768 VL. Guantity demanded x, = 73 poThe commodhly demanded for face le when price=o is given by -4x+1060=o - 426 = -1060 Is 1000 = 265 The quantity demanded free